QUESTION 1 Find the coordinate vector [p(x)]ℬ of p(x) = 5 − x + 3x 2 with respect to the basis ℬ = {u1, u2, u3 } of P2 where u1 = 1 − x + 3x 2 , u2 = 2 and u3 = 3 + x 2


1
Expert's answer
2022-04-19T02:28:01-0400

"In\\,\\,basis\\,\\,\\left\\{ 1,t,t^2 \\right\\} \\\\p=\\left[ \\begin{array}{c}\t5\\\\\t-1\\\\\t3\\\\\\end{array} \\right] \\\\Transition\\,\\,matrix\\\\T=\\left[ \\begin{matrix}\t1&\t\t2&\t\t3\\\\\t-1&\t\t0&\t\t0\\\\\t3&\t\t0&\t\t1\\\\\\end{matrix} \\right] \\\\The\\,\\,coordinates\\\\p'=T^{-1}p=\\left[ \\begin{matrix}\t1&\t\t2&\t\t3\\\\\t-1&\t\t0&\t\t0\\\\\t3&\t\t0&\t\t1\\\\\\end{matrix} \\right] ^{-1}\\left[ \\begin{array}{c}\t5\\\\\t-1\\\\\t3\\\\\\end{array} \\right] =\\left[ \\begin{matrix}\t0&\t\t-1&\t\t0\\\\\t0.5&\t\t-4&\t\t-1.5\\\\\t0&\t\t3&\t\t1\\\\\\end{matrix} \\right] \\left[ \\begin{array}{c}\t5\\\\\t-1\\\\\t3\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t1\\\\\t2\\\\\t0\\\\\\end{array} \\right] \\\\"


I find the inverse matrix with Gauss method: